# Quantum mechanics provides the physical basis of teleological evolutions

### v4

@@ Abstract
We show that the quantum computational speedup is due to the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living for which the teleological notion was originally conceived.
@@ 1 Introduction
Let us briefly retrace the two and a half millennia history of the teleological notion: attraction exercised by a final cause. Plato and Aristotle applied it the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution. With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy. It was eventually dismissed by science (it was classical science) for the alleged absence of a physical basis, even ascribed, for its finalistic character, to the category of beliefs. All this was before the advent of quantum mechanics, but has remained almost unchanged with it. Today, the situation of the teleological notion is somewhat complex. However, to the best of our knowledge, outside the present work, in physics, there are no researches on it. Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science. This is because quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Since showing this will be simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
We, with coauthors, already went against this position in the evolutionary, retrocausal, approach to the quantum computational speedup [1, 2, 3, 4, 5, 6]. In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obeyed the following rule:
@@
We called the rule in question the advanced knowledge rule. Note that knowing in advance one of the possible halves of the information about the solution one will produce and measure in the future and benefiting from this knowledge to produce the solution with fewer computation steps is a causal loop. We can therefore say that, according to this rule, the quantum speedup is explained by a mutually exclusive or of causal loops, one for each possible way of taking half of the information about the solution.
By the way, here and in the following, by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible. Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question. Also the notion of finalism acquires a precise definition. The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem). We will see further on that the physical definition of teleological evolution must be that of an evolution between an initial and a final measurement outcome.
Until now, this advanced knowledge rule was, so to speak, an empirical rule shown to be satisfied by all major quantum algorithms. In the present work, we should have demonstrated it, incidentally, in a surprisingly simple way. Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
As a matter of fact, we will provide two redundant demonstrations of this rule, which should not be despised given its destabilizing character. Let us outline them.
One is an application of classical logic to what the quantum algorithms does: providing a quantum speedup with respect to what is classically possible.
By the way, we will have to use a quantum description of the quantum algorithms that is not the customary one. The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement.
We complete this customary quantum description by adding, to the description of the process of solving the problem, that of the upstream process of setting the problem. The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow. Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic.
Having said this, the outline of the first demonstration can be very short. As we shall see in the next section, under the present more complete quantum description of the quantum algorithm, the classical logic description of the quantum speedup turns out to be indeed the advanced knowledge rule.
Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete. Indeed, it is simply completed by time-symmetrizing it. We shall see that the time-symmetrization of the more complete quantum description of the quantum algorithm generates the quantum superposition of the causal loops involved in the advanced knowledge rule.
The second demonstration of the advanced knowledge rule simply comes from the fact that the time-symmetrization of the more complete quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far. As we shall see, it is to satisfy an atemporal symmetry implicit in this description. As we have seen, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—we have thus a demonstration of the rule.
Naturally, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
A further step, important for the possibility of extending the teleological character of quantum algorithms to the evolutions of the living, is the following generalization.
For simplicity, let us limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty. We can then say that the quantum algorithm unitarily builds up one-to-one correlation between two initially unrelated observables—the problem-setting and the solution of the problem—without ever changing the eigenvalue of the observable selected by the initial measurement—the problem-setting.
We shall show that any building up of quantum correlation of this kind, also outside quantum computation, can occur with quantum speedup and therefore be teleological in character. Note that the “can occur” means that there is certainly the room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation. By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
We go now to the evolutions of the living for which the teleological notion was conceived.
First, let us note that the dismissal of the teleological notion was not without (scientific) regret; indeed, it meant doing without a possibly enlightening relation between the notion of evolution of life and that of the functional structure of the living. We cite two authoritative objections to this dismissal. One is a famous sentence by the evolutionary biologist Haldane [8]:
The other is the position of the natural philosopher Nagel [9], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view. In our possibly biased opinion, this would be a case in which common sense prevails over a de facto no-retrocausality precept that, according to the present work, would be a prejudice.
Let us now go to what the teleological character of quantum algorithms can say to the evolutions of the living.
Naturally, the first observation must be on the enormous difference between quantum algorithms and the evolutions of the living. The former have maximum durations of a few dozen microseconds and a number of quantum bits limited to the order of one thousand; beyond that, in today’s implementations, there is quantum decoherence. The latter have a duration of billions of years and can have the size of the Earth. However, the very context we are dealing with suggests a way out. The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11]. Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) could acquire the form of a quantum algorithm and, as such, it could occur with quantum speedup and thus be teleological in character.
Let us describe how things should go.
The Fine-tuned universe version of the Anthropic Principle is about the extraordinary theoretical fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
Under the quantum cosmological assumption, this version of the Principle can be formalized as follows.
We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
Of course, at the beginning of time, these two observables must be unrelated: no matter the value of the fundamental constant, the value of life is “absence of life”. At the time of today, the unitary evolution of the universe toward the value of life has built up one-to-one quantum correlation between the value of the fundamental constants since the beginning of time and the value of life, of course, without ever changing the former value.
We can therefore say that the evolution of the universe toward the value of life can occur with quantum speedup and thus be teleological in character; the “can occur” is in the sense that there is certainly the room for its occurring with quantum speedup. Of course, whether the quantum speedup is achieved depends on the character of the unitary evolution. For what was said about quantum random search algorithms, we might think that, if the search for the value of life by the unitary evolution of the universe is random, it should occur with a quantum speedup.
In the universe we are in, the evolution of the universe toward the value of life becomes evolution of the universe toward life; of course, all the above also applies to the latter evolution. We shall see that, as a consequence, the evolution of life on Earth, in principle the trace over all the rest of the evolution of the universe toward life, would inherit the possible teleological character of the evolution of the universe toward life. For consistency, the same would apply to the evolution of species.
Note that the fact that a teleological evolution is also an evolution endowed with a quantum speedup is a novelty for the teleological notion. It could contribute to explaining steps of natural evolutions that are too rapid to be explained by the progressive character of Darwinian evolutions.
Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption. Anyway, all this should be quite different from the present widespread belief that the notion of a teleological evolution of the living is an absurdity. At least, it should cast a doubt on this position of science.
Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they would acquire a physical basis. Of course, this would not touch on the immense practical value of the mechanistic and reductionist view. That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe holds only in principle: obviously, in practice, we cannot have such descriptions. However, this contrast between the quantum and the classical view might pinpoint a direction of the research. At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system (standing for the evolution of life on Earth) in the longer term with a classical character in the shorter term. This might even be a further perspective under which to study the possibility of unifying quantum and classical mechanics.
2 Time-symmetric explanation of the quantum speedup
We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality. Since this is a way of thinking quantum computation that is radically different from the usual one, we provide it first at a conceptual level. All formalizations are postponed in the Appendix. The section is divided into subsections corresponding to the various arguments.
Let us first acknowledge the fact that that of the quantum computational speedup has been one of the fundamental discoveries of the past century. Together with that of quantum communication, it has created an entirely new branch of physics. However, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
2.1 On the explanation of the quantum computational speedup
We repeat here for convenience the advanced knowledge rule.
@@ 2.1 On the explanation of the quantum computational speedup
Of course, this rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information. Let us discuss this point. To the best of our knowledge, the absence of a unified quantitative explanation has never been acknowledged by the scientific community.
@@
People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition. Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, cannot be considered its explanation. By the way, what distinguishes science from opinion is, in particular, its quantitative character. The fact that the most popular explanation of the quantum speedup, a quantitative feature, is qualitative in character, would seem to be rather unscientific. The only justification we can see is that this is a way of removing the fact that, forty years after its discovery—an explanation of the quantum speedup is still missing. This is naturally our opinion. However, as a matter of fact, outside our (unconventional) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
Of course, as already noted, if the only way of providing the unified quantitative explanation in question is by resorting to retrocausality and finalism, the de facto rejection of these notions by the scientific community might have played a role.
@@
Neither can be considered an explanation the fact that each quantum algorithm naturally explains its quantum speedup. In fact, it would be an explanation completely different for different quantum algorithms. It would be like having different explanations for different throws of a stone without knowing that they all describe a parabola. Obviously, unification is a fundamental objective of any description of physical phenomena.
Summing up, in our judgment, for what concerns the current mainstream explanation of the quantum speedup, it is the case of saying that the emperor is naked.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it has not been demonstrated; second, it goes against the de facto ban on retrocausality. In this work, we should have demonstrated it, moreover, in a surprisingly simple way. As a matter of fact, we shall be dealing with two independent demonstrations, which is not to be despised because of the destabilizing character of the rule.
2.2 First demonstration of the advanced knoledge rule
@@
We have already provided in the Introduction the quantum description of quantum algorithms extended to the process of solving the problem. An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^1] The problem-setter could unitarily change it at will; this, here, would change nothing.. The unitary computation and the final measurement of the corresponding solution follow. For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
@@
Now, we should note that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. To present ends, we should relativize it with respect to the problem-solver who, for the fact of computing the solution of the problem, is part of the process. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the completed quantum description of the quantum algorithm with respect to her. Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would tell her the solution of the problem before she begins her problem-solving action. Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easy. It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of her unitary computation of the solution; see Section (2) of the Appendix.
@@
We can go now to the first demonstration of the advanced knowledge rule. It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
By the way, involving only the measurements of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [12]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup. (ii) In Long’s [13] version, it always provides the solution with certainty; this simplifies things. (iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iv) It is optimal in character; this, of course, constitutes an important reference. By the way, note that the notion of searching for something and finding it is most universal and that this universality is completely represented by Grover algorithm.
We show that the classical logic description of the quantum speedup delivered by Grover algorithm very directly implies the advanced knowledge rule. We begin with the simplest instance of Grover algorithm. As we shall soon see, everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
The problem to be solved is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers. Alice, the problem-solver, is to locate the ball by opening drawers. In the worst classical case, she has to open three drawers. If the ball is only in the third drawer opened, she has solved the problem. If not, it must be in the only drawer not yet opened, and she has solved the problem of locating the ball as well.
@@
There is a classical logic equivalent to the fact that the quantum algorithm always solves the problem by opening a single drawer. By the way, since the fact that a single drawer is opened is logically true in itself, we can forget the specification that this is done for a quantum superposition of the four drawer numbers.
The logical equivalence is as follows. Logically, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers (one of which, of course, the one with the ball) and took advantage of this knowledge to solve the problem by opening just one of them. In fact, if the ball is in the opened drawer, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. This also means knowing in advance half of the information that specifies the solution of the problem that will be produced and measured in the future.
@@
Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way, one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer. Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, for her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that the information in question retrocausally comes to her from her future measurement of the solution. This is formalized in Section (3) of the Appendix. However, before looking at it, one should read the description of the time-symmetrization of the quantum algorithm that we shall provide in a moment.
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future (that the ball is in two rather than four drawers) and can use this knowledge to produce the solution by opening just one drawer. There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup. In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.

### v7

@@ Abstract
We show that the quantum computational speedup of quantum algorithms is due to their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living for which the teleological notion was originally conceived.
@@ 1 Introduction
Once applied to the evolutions of the living by natural philosophy, the notion of teleological evolution—of an evolution attracted by a final goal—was dismissed with the advent of modern science for the alleged lack of a physical basis. For its finalistic character, it was also ascribed to the category of beliefs. In this work, we shall show that the relatively recent discovery of the quantum computational speedup, the fact that quantum algorithms can solve problems with fewer logical operations (computation steps) than the minimum classically possible, should change this position of science. Indeed, we shall show that the quantum speedup of quantum algorithms is due to their teleological character.
In previous works of coauthors and ours [1, 2, 3, 4, 5, 6], we developed an unconventional, time-symmetric approach to the quantum speedup. In particular, in [4], we showed that:
In quantum algorithms, it is as if the problem-solver, before beginning her problem-solving action, knew in advance any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to compute the solution with fewer computation steps; the number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way benefiting from the advanced knowledge in question.
We have called the above the advanced knowledge rule. In the above publications, it was given as an empirical rule holding for all the quantum algorithms examined, which comprised the major ones. In the present work, we should have demonstrated it. Note that this rule, in addition to providing a unified quantitative explanation of all quantum speedups that is absent in mainstream quantum information, implies the teleological character of quantum algorithms. Indeed, according to it, they are evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce and measure in the future (the solution of the problem again): Without parentheses and their contents, this can well be considered the definition of teleological evolution. This also clarifies the finalistic character of the notion. A teleological evolution is not attracted by some superior scientifically undefinable goal, as argued by the opponents of the notion, but, quite simply, by its own end product.
Note that the advanced knowledge rule establishes the number of computation steps needed to solve an (oracle) problem in an optimal quantum way. Whether the quantum algorithm is optimal is another thing. We can say that the advanced knowledge rule tells us the room for the quantum speedup.
Of course, the advanced knowledge rule, its necessarily implying a form of retrocausality, is a destabilizing notion. Let us address in the first place this critical aspect of the work.
Knowing in advance, on the part of the problem-solver, any of the possible halves of the information about the solution she will produce and measure in the future and using this knowledge to produce the solution with fewer computation steps is obviously a causal loop. As there is a plurality of ways of taking half of the information about the solution, we can say that, according to the advanced knowledge rule, the quantum speedup is explained by the mutually exclusive or of these causal loops.
As far as we know, causal loops are completely absent from ordinary non-relativistic quantum mechanics. Of course, the current quantum description of quantum algorithms is causal in character—i.e. with causality always going in the same time-direction. Naturally, also this quantum description explains the quantum speedup of each quantum algorithm, although in a different way for each quantum algorithm—we shall go back to this further below.
We have to justify the discrepancy between what is described by the advanced knowledge rule and the current quantum description of quantum algorithms. It turns out that the current quantum description is incomplete and, once completed, it comprises the advanced knowledge rule. It is incomplete in a double way; the first is trivial (but should be amended first), the second is fundamental.
The trivial incompleteness consists in the fact that the quantum description is limited to the unitary computation of the solution where the quantum speedup resides and the final measurement of the latter. Of course, the canonically complete quantum description of a quantum process consists of initial measurement, unitary evolution, and final measurement. This incompleteness is also trivially amended. To the quantum description of the process of solving the problem, we should add that of the upstream process of setting the problem. An initial measurement of the problem-setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and final measurement of the corresponding solution follow [^1] We exemplify what we mean by problem-setting. Consider the following problem: Bob, the problem-setter, hides a ball in a chest of drawers, Alice, the problem-solver, should locate it by opening drawers. The number of the drawer with the ball selected by Bob is the problem-setting.. We call this the ordinary quantum description of the quantum algorithm; “ordinary” because it is still the causal quantum description of mainstream science.
This ordinary quantum description is still incomplete, this time in the fundamental way. Further below, we shall see that it is exactly the same incompleteness lamented by Einstein et al in the case of the quantum description of quantum nonlocality [7].
Before proceeding, let us simplify the situation by requiring that the solution of the problem is an invertible function of the problem-setting and is always determined with certainty (of course, the problem-setting and the solution are two commuting observables). We should also keep in mind that the unitary computation of the solution never changes the problem-setting.
With the above, we can go to the amendment of the fundamental incompleteness of the quantum description we are dealing with. It turns out that this is done by time-symmetrizing it. It is a time-symmetrization inspired to time-symmetric quantum mechanics [8, 9, 10]. As we shall see, the time-symmetrization in question also turns out to be mandatory to satisfy an atemporal symmetry implicit in quantum algorithms and highlighted by their present (ordinary) quantum description.
Things go as follows. Ordinarily, the information that specifies the problem-setting and the corresponding solution is all selected by the initial measurement of the problem-setting. We should instead require that this selection evenly shares between the initial measurement of the problem-setting and the final measurement of the solution. As there is a plurality of ways of evenly sharing, we should take their quantum superposition.
In each time-symmetrization instance, the outcome of the initial measurement should propagate forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. Time-symmetrically, the outcome of the final measurement should propagate backward in time, by the inverse of the unitary computation of the solution, until replacing the previous outcome of the initial measurement. As we shall see, this backward in time propagation tells the problem solver, before she begins her problem-solving action, one of the possible halves of the information about the solution. Indeed, the selection performed by the final measurement affects the initial measurement outcome, changing the initial complete ignorance of the solution on the part of the problem-solver into knowledge of one of the possible halves of the information about it [^2] On the possibility that a future selection affects a past measurement outcome see [10].. She can then use this knowledge to produce the solution with fewer computation steps. In other words, this time-symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is implicit in the advanced knowledge rule. Besides reconciling the ordinary quantum description of the quantum algorithms with the advanced knowledge rule, this is the demonstration of the latter—we should keep in mind that the time-symmetrization we are dealing with is mandatory.
By the way, we shall provide another, independent, demonstration of the advanced knowledge rule. The fact that the ordinary quantum description of the quantum algorithm involves only the measurements of commuting observables (the problem-setting and the solution of the problem) makes it subject to classical logic. Now, as we shall see, the classical logic description of the fact that the quantum algorithm computes the solution with fewer computation steps than the minimum possible in the classical case very directly implies the advanced knowledge rule. It is interesting to provide also this demonstration because it very clearly highlights the interplay between time-forward and time-backward causality in unitary quantum evolutions that produces the quantum speedup.
Let us make another specification in view of what will follow. Of course, the sending of information backward in time operated by the time-symmetrization of the ordinary quantum description of the quantum algorithm must be along a time-reversible process. Here, it is the unitary computation of the solution. The special character of its time-reversibility should be underlined. Unlike the classical time-revesibility, which is necessarily an idealization, the present kind is both ideal and objective. This is, of course, an essentially quantum feature that stands at the basis of the present time-symmetric explanation of the quantum speedup.
We have seen that demonstrating the advanced knowledge rule also amounts to demonstrating the teleological character of quantum algorithms. Of course, this would already falsify the current opinion that teleological evolutions cannot have a physical basis. However, their present physical basis is necessarily in the quantum world and the teleological notion was applied to the evolutions of the living. We shall show that, under the quantum cosmological assumption and for the Fine-tuned Version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms identically applies to the evolutions of the living. Under the quantum cosmological assumption, the evolution of the universe toward life becomes quantum. For that version of the Anthropic principle, it acquires the form of a quantum algorithm and, as such, can be endowed with quantum speed up and consequently be teleological in character. The evolution of life on Earth, in principle the trace over all the rest of the former evolution, and, with it, the evolutions of life, would inherit that possible teleological character.
In summary, the time-symmetrization we are dealing with (and the corresponding completion of the quantum description) turns out to have a special unification power. Besides providing a unified quantitative explanation of all quantum speedups (the one provided by the advanced knowledge rule), it unifies the explanation of the quantum speedups with that of quantum nonlocality and, under the quantum cosmological assumption, it unifies the teleological character of quantum algorithms with that of the (macroscopic) evolutions of the living for which the teleological notion was originally conceived.
2 Time-symmetric explanation of the quantum speedup
We provide a description of quantum algorithms and their quantum speedups open to the time-symmetric coexistence of causality and retrocausality along unitary quantum evolutions. In this section, we provide it at a conceptual level, but making reference to the formalizations postponed in an Appendix. The section is divided into subsections corresponding to the various arguments.
2.1 Critique of the current way of viewing quantum computation
Let us first acknowledge that that of the quantum computational speedup has been one of the fundamental discoveries of the past century. Together with that of quantum communication, it has created an entirely new branch of physics. However, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation, likely related to the rejection of the notion of retrocausality.
We repeat here for convenience the advanced knowledge rule; of course, it plays an essential role in the present work.
@@ 2.1 Critique of the current way of viewing quantum computation
This rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information. Let us discuss this point. To the best of our knowledge, this absence has never been acknowledged by the scientific community.
@@
People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition. Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, it cannot be considered its explanation. Of course, science essentially differs from opinion because of its quantitative character; then, settling for a qualitative explanation of a quantitative feature is rather unusual. This is our opinion, of course. However, as a matter of fact, outside the present retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
@@
Neither can be considered a satisfactory explanation of the quantum speedup the fact that each quantum algorithm naturally explains its quantum speedup. In fact, it would be an explanation completely different for different quantum algorithms. It would be like having different explanations for different throws of a stone without knowing that they all describe a parabola. Naturally, unification of explanations is a fundamental objective of science.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. In this section, we shall provide the two demonstrations of it, beginning with the one relying on the classical logic description of the quantum speedup.
2.2 First demonstration of the advanced knowledge rule
Let us recall the quantum description of quantum algorithms extended to the process of solving the problem already provided in the Introduction:
@@ 2.2 First demonstration of the advanced knowledge rule
An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^3] The problem-setter could unitarily change it at will; this, here, would change nothing.. The unitary computation and the final measurement of the corresponding solution follow.
For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty. Table I of Section (1) of the Appendix provides the formalization of this extended quantum description in the simplest instance of Grover algorithm. Here, the problem is as follows: Bob, the problem-setter, hides a ball in a chest of four drawers, Alice, the problem-solver, is to locate it (i.e. identify the number of the drawer with the ball) by opening drawers.
@@
Note that the quantum description of the quantum algorithm of Table I is with respect to the customary observer external to the quantum process. To present ends, we should relativize it with respect to the problem-solver who, for the fact of computing the solution of the problem, is part of the process. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the (completed) quantum description of the quantum algorithm with respect to her. Now, the problem-setting selected by the initial measurement should be concealed from her: as we shall see, it would tell her the solution of the problem before she begins her problem-solving action: Of course, at this time, instead, she is completely ignorant of the solution. Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easy. It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of her unitary computation of the solution; see Table II of Section (2) of the Appendix.
@@
We can now go to the first demonstration of the advanced knowledge rule. It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that are less than the minimum possible in the classical case.
By the way, since their initial and final measurements are of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
Now, we have to choose the quantum algorithm to work with. We need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [13]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) In Long’s version [14], it always provides the solution with certainty; this simplifies things. (ii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iii) It is optimal in character; this, of course, constitutes an important reference. (iv) It is a most universal quantum algorithm: it can easily be adapted to solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup.
Let us show that (under classical logic) the quantum speedup over what is classically possible delivered by Grover algorithm logically implies the advanced knowledge rule. We begin with the simplest instance of this quantum algorithm: everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
The problem to be solved is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers. Alice, the problem-solver, is to locate the ball by opening drawers. In the worst classical case, she has to open three drawers: If the ball is only in the third drawer opened, she has solved the problem. If not, it must be in the only drawer not yet opened, and she has solved the problem of locating the ball as well.
@@
There is a classical logic equivalent of the fact that the quantum algorithm always solves the problem by opening a single drawer. By the way, since the fact that a single drawer is opened is logically true in itself, we can forget the specification that this is done for a quantum superposition of the four drawer numbers.
The logical equivalence is as follows. Logically, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers (in a mutually exclusive way one of the three pairs of drawers in which the ball must be) and took advantage of this knowledge to solve the problem by opening just one of them. In fact, if the ball is in the opened drawer, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. By the way, this also means knowing in advance half of the information that specifies the solution of the problem that will be produced and measured in the future.
@@
Always knowing in advance that the ball is in a given pair of drawers is logically necessary to always solve the problem by opening just one drawer. Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, for her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that the information in question retrocausally comes to her from her future measurement of the solution. This is formalized in Section (3) of the Appendix. However, before looking at it, one should read the description of the time-symmetrization of the quantum algorithm that we shall provide in a moment.
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future (that the ball is in two rather than four drawers) and can use this knowledge to produce the solution by opening just one drawer. There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup.